Find the range the standard deviation and the variance for the given samples

1
Expert's answer
2021-12-15T16:52:12-0500

Assume the data is: 48, 91, 87, 93, 59, 68, 92, 100, 81

Solution:

"48, 59, 68, 81,87,91, 92, 93, 100"

"Range=100-48=52"

"mean=\bar{x}=\dfrac{\displaystyle\sum_{i=1}^nx_i}{n}=\dfrac{1}{11}(48+59+68+81+87"

"+91+92+93+100)=\dfrac{719}{9}"

"\approx79.9"

"Variance=s^2=\dfrac{\displaystyle\sum_{i=1}^n(x_i-\bar{x})^2}{n-1}"

"=\dfrac{1}{8}((48-\dfrac{719}{9})^2+(59-\dfrac{719}{9})^2+(68-\dfrac{719}{9})^2"

"+(81-\dfrac{719}{9})^2+(87-\dfrac{719}{9})^2+(91-\dfrac{719}{9})^2"

"+(92-\dfrac{719}{9})^2+(93-\dfrac{719}{9})^2+(100-\dfrac{719}{9})^2)"

"\approx311.6"

"s=\sqrt{s^2}\approx17.7"